# Factor copula models for non-Gaussian longitudinal data

### v2

@@ Abstract
This article presents factor copula approaches to model temporal dependency of non-Gaussian (continuous/discrete) longitudinal data. Factor copula models are canonical vine copulas which explain the underlying dependence structure of a multivariate data through latent variables, and therefore can be easily interpreted and implemented to unbalanced longitudinal data. We develop regression models for continuous, binary and ordinal longitudinal data including covariates, by using factor copula constructions with subject-specific latent variables. Considering homogeneous within-subject dependence, our proposed models allow for feasible parametric inference in moderate to high dimensional situations, using two-stage (IFM) estimation method. We assess the finite sample performance of the proposed models with extensive simulation studies. In the empirical analysis, the proposed models are applied for analysing different longitudinal responses of two real world data sets. Moreover, we compare the performances of these models with some widely used random effect models using standard model selection techniques and find substantial improvements. Our studies suggest that factor copula models can be good alternatives to random effect models and can provide better insights to temporal dependency of longitudinal data of arbitrary nature.
Correspondence: Subhajit Chattopadhyay, email: subhajitc.stat.isi@gmail.com
Keywords: factor copula; temporal dependency; GLM; latent variable model; longitudinal data.
@@ 1 Introduction
Longitudinal studies consist of collecting repeated measurements from subjects over time with an aim to understand the relationship with some explanatory variables (covariates) as well as the dependence mechanisms among these measurements. These measurements can be of different nature (discrete/continuous) depending upon the investigation. The statistical modeling of continuous longitudinal data have been widely studied due to availability of various flexible multivariate distributions, but for the discrete cases the approaches are rather limited. Verbeke & Molenberghs (1997), Davis et al. (2002) or Fitzmaurice et al. (2008) have prevalent overviews and concepts of longitudinal data analysis. To analyse continuous longitudinal data, the most popular class of models are the linear mixed models (LMM) which are based on multivariate normality assumption of the repeated measurements (Laird & Ware (1982)). Several other approaches have been proposed to relax distributional assumptions to handle non-Gaussian longitudinal data. One of which among these, getting increasingly popular in recent times is the copula framework.
@@
One of the earlier applications of copulas to longitudinal data analysis was by Lambert & Vandenhende (2002). Sun et al. (2008) proposed the use of multivariate elliptical copulas to model longitudinal data. Shi et al. (2016) considered unbalanced longitudinal study design using Gaussian copula. In recent times, Killiches & Czado (2018) adopted a flexible D-vine copula approach to model unbalanced continuous longitudinal data by using linear mixed models as marginals. D-vine copulas offer a great flexibility and imply a natural temporal ordering of the repeated measurements, but can be computationally demanding in high-dimension due to involvement of many parameters. Panagiotelis et al. (2012) used pair-copulas to model dependence of discrete data without involving covariates. Zilko & Kurowicka (2016) employed pair-copulas to model mixed data. Sefidi et al. (2021) considered D-vine copulas to construct various multivariate distributions using power series marginal distributions, but their applications were limited to dimension $d=3$. Factor copulas introduced by Krupskii & Joe (2013), provide with similar level of flexibility but with fewer dependence parameters. In factor copula models, dependence between the observed variables are explained through some latent variables. Nikoloulopoulos & Joe (2015) extended the factor copula approach to model discrete valued item response data. Kadhem & Nikoloulopoulos (2021) implemented some factor copula models in mixed response type social science data.
@@
Factor copula models are beneficial in the sense that these can be applied to response variables of any nature without having numerical complications in moderate to high dimensions. These are truncated vine-copula models which involves both observed and latent variables. Based on the selection of bivariate linking copulas, factor copulas provide a wide range of asymmetric, tail and nonlinear dependence to the data. Kadhem & Nikoloulopoulos (2021) emphasized that factor copula models are more interpretable and fit better than vine copula models and these are closed under marginals, that is, lower-order marginals belong to the same parametric family of copulas and a different permutation of the observed variables has exactly the same distribution. In this article our main contribution is to introduce factor copulas to model the temporal dependency of unbalanced non-Gaussian longitudinal data. For continuous responses we considered generalized linear models (GLM) for the marginals and for binary or ordinal responses we link factor copulas with an underlying normally distributed latent variable. This supports us to use subject specific covariates, which are the key components in longitudinal data analysis. We restrict with univariate analysis of the longitudinal responses of different nature. Considering missing data mechanism to be ignorable, we estimate the model parameters using IFM method (inference function of margins). Factor copula models can also be compared to random effect models such as generalized linear mixed models (GLMM) since they both involve latent variables on the subject level. Another contribution of this article is to compare the performances of factor copula models with random effect models, though they could have different interpretations under different circumstances.
@@
The article is organized as follows: in Section 2, we describe factor copula constructions for (i) continuous responses with generalized linear models and (ii) discrete responses with latent variables. In Section 3 we describe the estimation techniques and the computational details. Section 4 describes the competing random effect models and some commonly used model selection techniques. In Section 5, we present diagnostics of the residuals of fitted factor copula models. Section 6 presents some extensive simulation studies to evaluate the finite sample performances of the proposed longitudinal models under some parsimonious specifications of factor copulas. We describe two motivating real world data sets in Section 7 and present our analysis. Section 8 concludes this article and describes some future extensions of the proposed models.
@@ 2 The factor copula models for continuous and discrete responses
It is well known that for longitudinal data, the repeated measurements exhibit a clear temporal ordering, hence for moderate to high-dimensional situations the use of factor copulas can be quite effective in terms of parametric inference. We assume that for the $i$-th subject the responses $\mathbf{Y}_{i}=(Y_{i1},\dots,Y_{in_{i}})^{\intercal}$ are conditionally independent given $p$ latent variables $V_{i1},\dots,V_{ip}$ which are independent and identically distributed (i.i.d) $U(0,1)$ (Krupskii & Joe (2013)). That is, a factor copula model is a conditional independence model which explains the dependence among the repeated measurements through some latent variables. The responses can be either continuous or discrete in nature. Here we describe the constructions of $1$-factor and $2$-factor copula models in the context of longitudinal data, but these can be extended up to $p(\geq 3)$ factors.
@@
where $\mathbf{\theta}^{*}_{i}=(\mathbf{\theta}_{i},\mathbf{\phi}_{i})^{\intercal}$ contains all the marginal and dependence parameters for the $i$-th subject. Thus in a $1$-factor copula model $n_{i}$ bivariate copulas couple each observed variable to the first latent variable. For a $2$-factor copula model we consider $2$ latent variables $V_{i1}$ and $V_{i2}$, i.i.d $\sim U(0,1)$. The joint distribution of $Y_{ij}$ and $V_{i2}$ given $V_{i1}$, can be written in terms of a copula $C_{j,2}(.|\delta_{ij})$ as
@@
$f_{n_{i}}(y_{i1},\dots,y_{in_{i}}|\mathbf{\theta}^{*}_{i})=\int_{0}^{1}\int_{0}^{1}\prod_{j=1}^{n_{i}}P(Y_{ij}=y_{ij}|V_{i1}=v_{i1},V_{i2}=v_{i2})dv_{i1}dv_{i2}\\ =\int_{0}^{1}\int_{0}^{1}\prod_{j=1}^{n_{i}}\Big{[}C_{j|1,2}(F_{ij|1}(\gamma(y_{ij})|v_{i1},\theta_{ij},\phi_{ij})|v_{i2},\delta_{ij})\\ -C_{j|1,2}(F_{ij|1}(\gamma(y_{ij}-1)|v_{i1},\theta_{ij},\phi_{ij})|v_{i2},\delta_{ij})\Big{]}dv_{i1}dv_{i2}.$ (2.13)
Here we are making the so called simplifying assumption that the conditional copula for the univariate distributions $F_{ij|1}$ and $v_{i2}$ does not depend on $v_{i1}$. The choice of bivariate linking copulas in $1$-factor and $2$-factor copula models is completely arbitrary, but in this article we restrict to the elliptical copulas in the sense that the dependence parameter have direct interpretation same as their bivariate distributions. We consider the bivariate Gaussian copula as
@@
where $T^{-1}$ is the quantile of the univariate Student-$t$ with $\nu$ degrees of freedom, and $T_{2}$ is the cdf of a bivariate Student-$t$ distribution with $\nu$ degrees of freedom and correlation parameter $\rho$. This additional degrees of freedom parameter $\nu$, accounts for possible tail dependence in the data. Small value of $\nu$, such as $2\leq 5$, leads to a model with more probabilities in the joint upper and joint lower tails compared with the normal copula.
@@ 3 Parameter estimation
The two most common frequentist technique of parameter estimation of copula based models is (i) maximum likelihood estimation (MLE), and (ii) inference function of margins (IFM) (Joe & Xu (1996) and Joe (2005)). The former is known to be the most efficient but can be computationally demanding for most the dependence models such as factor copula models. Here we implement the IFM method in which the model parameters are estimated in two separate stages. First we estimate all marginal parameters assuming independence, by maximizing pseudo likelihood of the form:
@@
and then using the parameter estimates from (3.1), we compute the uniform samples: $u_{ij}=F(y_{ij}|\hat{\theta}_{ij})$ for the continuous case, and $u_{ij}=F(\gamma(y_{ij})|\hat{\theta}_{ij})$, $u_{ij}^{-}=F(\gamma(y_{ij}-1)|\hat{\theta}_{ij})$ for the discrete case, respectively where $i=1,\dots,m,j=1,\dots,n_{i}$. Thereafter we estimate the parameters of $1$-factor copula models for the continuous case by maximizing the pseudo likelihood:
@@
$l(\mathbf{\phi},\mathbf{\delta}|\mathbf{u})=\sum_{i=1}^{m}\log\int_{0}^{1}\int_{0}^{1}\prod_{j=1}^{n_{i}}\Big{[}C_{j|1,2}(C_{j|1}(u_{ij}|v_{i1},\phi_{ij})|v_{i2},\delta_{ij})\\ -C_{j|1,2}(C_{j|1}(u_{ij}^{-}|v_{i1},\phi_{ij})|v_{i2},\delta_{ij})\Big{]}dv_{i1}dv_{i2}.$ (3.7)
@@
where $D(\hat{\mathbf{\theta}}^{*})$ is a block diagonal matrix and $M(\hat{\mathbf{\theta}}^{*})$ is a symmetric positive definite matrix. More details regarding estimation of the matrix can be found in Joe & Xu (1996) or Joe (2014). To estimate the parameters we use optim function, and to estimate the information matrix associated with the parameter estimates we use numderiv function in R.
@@ 4 Comparison with random effect models
Random effects are generally used to introduce subject-specific effects to the linear predictor of the models as well as within-subject dependency or temporal dependency (see, Shi & Valdez (2014)). Factor copula models can be compared with the random effect models since both of these models involves unobservable latent variables. In this article we also consider random effect models for accounting the temporal dependency as competitors by extending the marginal models introduced in Section 2. Referring to (2.16), we consider generalized linear mixed models as
@@
$l(\mathbf{\theta}|\mathbf{y})=\sum_{i=1}^{m}\int\prod_{j=1}^{n_{i}}f(y_{ij}|\mathbf{b}_{i})h(\mathbf{b}_{i})d\mathbf{b}_{i},$ (4.3)
@@ 5 Residual analysis
Here we show Rosenblatt’s transformation (Rosenblatt (1952)) is readily applicable for factor copula models. Masarotto & Varin (2012) or Hofert & Mächler (2014) previously used this method to validate copula assumptions for multivariate models. This method transforms dependent random variables into independent uniform random variables in the unit interval. Consistent with the setting considered in Section 2 let
@@ 6 Simulation studies
To illustrate the performance of the estimation methods for our proposed class of models we undertake some simulation studies imitating the features of the data sets under consideration. We simulate data sets from the proposed factor copula based models and monitor the parametric inference. We consider two different sample sizes, $m=\{200,500\}$ and set the maximum number of longitudinal responses to $d=10$. Then to mimic the unbalanced nature of the data, for each unit we generate $n_{i}\sim Binomial(d,p=0.8)$, (for $i=1,\dots,m$), that is we prune the data set. Since the true parameters are known, we first sample $\mathbf{U}_{i}$, from $n_{i}$-variate ($n_{i}\leq d$) $1$-factor and $2$-factor copulas (for $i=1,\dots,m$). Thereafter, we use PIT on $\mathbf{U}_{i}$ to generate per-unit responses $\mathbf{Y}_{i}$. Since the dimension for each response is different, we assume same bivariate copulas for each factor with exchangeable parameters. This leads to reduction in the number of estimable parameters from the model. We generate the continuous responses from the following model -
@@
where $\epsilon_{ij}(i.i.d)\sim N(0,1)$. Here we assign the regression coefficients as, $\beta_{0}=-0.5,\beta_{1}=-0.5,\beta_{2}=0.2,\beta_{3}=0.2$. Finally we generate the ordinal responses from the following model -

### v3

@@ Abstract
This article introduces factor copula approaches for modeling the temporal dependence of non-Gaussian (continuous/discrete) longitudinal data. Factor copula models, which are canonical vine copulas, elucidate the underlying dependence structure of multivariate data through latent variables, thereby facilitating interpretation and implementation for unbalanced longitudinal data. We develop regression models for continuous, binary, and ordinal longitudinal data, incorporating covariates, using factor copula constructions with subject-specific latent variables. With consideration for homogeneous within-subject dependence, our proposed models enable feasible parametric inference in moderate to high dimensional scenarios, employing a two-stage (IFM) estimation method. We evaluate the finite sample performance of the proposed models through extensive simulation studies. In empirical analysis, we apply the proposed models to analyze various longitudinal responses from two real-world datasets. Furthermore, we compare the performance of these models with widely used random effect models using standard model selection techniques and observe substantial improvements. Our findings suggest that factor copula models can serve as viable alternatives to random effect models, offering deeper insights into the temporal dependence of longitudinal data across diverse contexts.
Keywords: factor copula; temporal dependence; GLM; latent variable model; longitudinal data.
@@ 1 Introduction
Longitudinal studies involve the systematic collection of repeated measurements from subjects over time, aimed at understanding their relationship with explanatory variables (covariates) and uncovering the underlying mechanisms of dependence among these measurements. Depending on the research focus, these measurements can manifest in various forms, ranging from discrete to continuous. While statistical methodologies for continuous longitudinal data have been extensively explored, benefiting from a rich array of flexible multivariate distributions, approaches for handling discrete cases remain relatively scarce. Seminal works by researchers such as Verbeke & Molenberghs (1997), Davis et al. (2002), and Fitzmaurice et al. (2008) offer invaluable insights and foundational concepts in longitudinal data analysis. Linear mixed models (LMM) have emerged as the go-to choice for analyzing continuous longitudinal data, resting on the assumption of multivariate normality for the repeated measurements (Laird & Ware (1982)). However, the landscape has seen the emergence of alternative methodologies aiming to relax stringent distributional assumptions and accommodate non-Gaussian longitudinal data. Among these, the copula framework has garnered increasing attention in recent years for its versatility and ability to capture complex dependencies within longitudinal data.
@@
The application of copulas to longitudinal data analysis traces back to Lambert & Vandenhende (2002), marking an early utilization of this methodology in this context. Sun et al. (2008) proposed the utilization of multivariate elliptical copulas to model longitudinal data, contributing to the diversification of copula applications. Shi et al. (2016) delved into unbalanced longitudinal study designs, employing Gaussian copulas to address the intricacies of such setups. More recently, Killiches & Czado (2018) adopted a flexible D-vine copula approach to model unbalanced continuous longitudinal data, leveraging linear mixed models as marginals. While D-vine copulas offer remarkable flexibility and naturally account for the temporal ordering of repeated measurements, they can pose computational challenges in high-dimensional settings due to the involvement of numerous parameters. Panagiotelis et al. (2012) explored the use of pair-copulas to model dependence in discrete data, devoid of covariate involvement, whereas Zilko & Kurowicka (2016) employed pair-copulas to address mixed data scenarios. Sefidi et al. (2021) extended the application of D-vine copulas to construct various multivariate distributions using power series marginal distributions, albeit within a limited dimensionality ($d=3$). Factor copulas, introduced by Krupskii & Joe (2013), offer comparable flexibility with fewer dependence parameters, explaining the interrelationships between observed variables through latent variables. Nikoloulopoulos & Joe (2015) extended the factor copula methodology to model discrete-valued item response data, while Kadhem & Nikoloulopoulos (2021) applied factor copula models to mixed response-type social science data, further broadening the scope of applications in this field.
@@
Factor copula models offer distinct advantages, particularly their applicability to response variables of any nature without encountering numerical complexities, even in moderate to high dimensions. These models, classified as truncated vine-copula models, incorporate both observed and latent variables. By selecting appropriate bivariate linking copulas, factor copulas afford a broad spectrum of asymmetric, tail, and nonlinear dependence structures to the data. Kadhem & Nikoloulopoulos (2021) underscored the interpretability and superior fit of factor copula models compared to vine copula models, noting their closure under marginals. This property ensures that lower-order marginals belong to the same parametric copula family, and different permutations of observed variables yield identical distributions. In this article, our primary contribution lies in introducing factor copulas as a tool for modeling the temporal dependence inherent in unbalanced non-Gaussian longitudinal data. For continuous responses, we employ generalized linear models (GLM) for the marginals, while for binary or ordinal responses, we link factor copulas with an underlying normally distributed latent variable. This framework enables the incorporation of subject-specific covariates, essential components in longitudinal data analysis. We focus on univariate analysis of longitudinal responses of varying natures, assuming responses are missing completely at random. Model parameters are estimated using the IFM method (inference function of margins). Factor copula models can be juxtaposed with random effect models, such as generalized linear mixed models (GLMMs), as both involve latent variables at the subject level. Another noteworthy aspect of our study is the comparison of the performances of factor copula models with random effect models, despite their potentially divergent interpretations in different contexts.
@@
The article is structured as follows: Section 2 details factor copula constructions for (i) continuous responses using generalized linear models and (ii) discrete responses employing latent variables. Section 3 elaborates on the estimation techniques and computational intricacies involved. Competing random effect models and commonly used model selection techniques are discussed in Section 4. In Section 5, diagnostics of the residuals of fitted factor copula models are presented. Extensive simulation studies to evaluate the finite sample performances of the proposed longitudinal models under parsimonious specifications of factor copulas are outlined in Section 6. Two motivating real-world datasets are described in Section 7, followed by the presentation of our analysis. Finally, Section 8 concludes the article and outlines potential future extensions of the proposed models.
@@ 2 The factor copula models for continuous and discrete responses
By the definition of longitudinal data, the repeated measurement tend to be serially correlated, hence for moderate to high-dimensional situations the use of factor copulas can be quite effective in terms of parametric inference. We assume that for the $i$-th subject the responses $\mathbf{Y}_{i}=(Y_{i1},\dots,Y_{in_{i}})^{\intercal}$ are conditionally independent given $p$ latent variables $V_{i1},\dots,V_{ip}$ which are independent and identically distributed (i.i.d) $U(0,1)$ (Krupskii & Joe (2013)). That is, a factor copula model is a conditional independence model which explains the dependence among the repeated measurements through some latent variables. The responses can be either continuous or discrete in nature. Here we describe the constructions of $1$-factor and $2$-factor copula models in the context of longitudinal data, but these can be extended up to $p(\geq 3)$ factors.
@@
where $\mathbf{\theta}^{*}_{i}=(\mathbf{\theta}_{i},\mathbf{\phi}_{i})^{\intercal}$ contains all the marginal and dependence parameters for the $i$-th subject and $c_{j,1}$ is the copula density function. Thus in a $1$-factor copula model $n_{i}$ bivariate copulas couple each observed variable to the first latent variable. For a $2$-factor copula model we consider $2$ latent variables $V_{i1}$ and $V_{i2}$, i.i.d $\sim U(0,1)$. The joint distribution of $Y_{ij}$ and $V_{i2}$ given $V_{i1}$, can be written in terms of a copula $C_{j,2}(.|\delta_{ij})$ as
@@
$f_{n_{i}}(y_{i1},\dots,y_{in_{i}}|\mathbf{\theta}^{*}_{i})=\int_{0}^{1}\int_{0}^{1}\prod_{j=1}^{n_{i}}P(Y_{ij}=y_{ij}|V_{i1}=v_{i1},V_{i2}=v_{i2})dv_{i1}dv_{i2}\\ =\int_{0}^{1}\int_{0}^{1}\prod_{j=1}^{n_{i}}\Big{[}C_{j|1,2}(F_{ij|1}(\gamma(y_{ij})|v_{i1},\theta_{ij},\phi_{ij})|v_{i2},\delta_{ij})-C_{j|1,2}(F_{ij|1}(\gamma(y_{ij}-1)|v_{i1},\theta_{ij},\phi_{ij})|v_{i2},\delta_{ij})\Big{]}dv_{i1}dv_{i2}.$ (2.13)
Here we are making the so called simplifying assumption that the conditional copula for the univariate distributions $F_{ij|1}$ and $v_{i2}$ does not depend on $v_{i1}$. The choice of bivariate linking copulas in $1$-factor and $2$-factor copula models is completely arbitrary, but in this article we restrict to the elliptical copulas in the sense that the dependence parameter have direct interpretation same as their bivariate distributions. This is also motivated from the distributional properties of elliptical copulas. We consider the bivariate Gaussian copula as
@@
where $T^{-1}$ is the quantile of the univariate Student-$t$ with $\nu$ degrees of freedom, and $T_{2}$ is the cdf of a bivariate Student-$t$ distribution with $\nu$ degrees of freedom and correlation parameter $\rho$. This additional degrees of freedom parameter $\nu$, accounts for possible tail dependence in the data. Small values of $\nu$, such as $2\leq 5$, leads to a model with more probabilities in the joint upper and joint lower tails compared with the normal copula.
@@ 3 Parameter estimation
The two most common frequentist technique of parameter estimation of copula based models is (i) maximum likelihood estimation (MLE), and (ii) inference function of margins (IFM) (Joe & Xu (1996) and Joe (2005)). The former is known to be the most efficient but can be computationally demanding for most the dependence models such as factor copula models. Here we implement the IFM method in which the model parameters are estimated in two separate stages. Note that ignorability property does applies for pseudo likelihood estimation (see, Molenberghs et al. (2011)). First we estimate all marginal parameters assuming independence, by maximizing pseudo likelihood of the form:
@@
and then using the parameter estimates from (3.1), we compute the uniform samples: $u_{ij}=F_{ij}(y_{ij}|\hat{\theta}_{ij})$ for the continuous case, and $u_{ij}=F_{ij}(\gamma(y_{ij})|\hat{\theta}_{ij})$, $u_{ij}^{-}=F_{ij}(\gamma(y_{ij}-1)|\hat{\theta}_{ij})$ for the discrete case, respectively where $i=1,\dots,m,j=1,\dots,n_{i}$. Thereafter we estimate the parameters of $1$-factor copula models for the continuous case by maximizing the pseudo likelihood:
@@
$l(\mathbf{\phi},\mathbf{\delta}|\mathbf{u})=\sum_{i=1}^{m}\log\int_{0}^{1}\int_{0}^{1}\prod_{j=1}^{n_{i}}\Big{[}C_{j|1,2}(C_{j|1}(u_{ij}|v_{i1},\phi_{ij})|v_{i2},\delta_{ij})-C_{j|1,2}(C_{j|1}(u_{ij}^{-}|v_{i1},\phi_{ij})|v_{i2},\delta_{ij})\Big{]}dv_{i1}dv_{i2}.$ (3.7)
@@
where $D(\hat{\mathbf{\theta}}^{*})$ is a block diagonal matrix and $M(\hat{\mathbf{\theta}}^{*})$ is a symmetric positive definite matrix. More details regarding estimation of the matrix can be found in Joe & Xu (1996) or Joe (2014). To estimate the parameters we use optim (Cortez & Cortez (2014)) function, and to estimate the information matrix associated with the parameter estimates we use numderiv (Gilbert et al. (2009)) function in R.
@@ 4 Comparison with random effect models
Random effects are generally used to introduce subject-specific effects to the linear predictor of the models as well as within-subject dependency or temporal dependence (see, Shi & Valdez (2014)). Factor copula models can be compared with the random effect models since both of these models involves unobservable latent variables. This means that models of this type are necessarily based on not fully verifiable assumption. For a detailed overview, we refer to the readers to see, Molenberghs et al. (2012) and Verbeke & Molenberghs (2010). In this article we also consider random effect models for accounting the temporal dependence as competitors by extending the marginal models introduced in Section 2. Referring to (2.16), we consider generalized linear mixed models as
@@
$l(\mathbf{\theta}|\mathbf{y},\mathbf{x})=\sum_{i=1}^{m}\int\prod_{j=1}^{n_{i}}f(y_{ij}|\mathbf{b}_{i})h(\mathbf{b}_{i})d\mathbf{b}_{i},$ (4.3)
@@ 5 Residual analysis
Here we show how Rosenblatt’s transformation (Rosenblatt (1952)) is readily applicable for factor copula models. Masarotto & Varin (2012) or Hofert & Mächler (2014) previously used this method to validate copula assumptions for multivariate models. This method transforms dependent random variables into independent uniform random variables in the unit interval. Therefore, can be used to validate the distributional assumption of different multivariate models. Consistent with the setting considered in Section 2 let
@@ 6 Simulation studies
To assess the performance of the estimation methods for our proposed models, we conduct simulation studies that emulate the characteristics of the datasets under consideration. We generate simulated datasets based on the proposed factor copula models and monitor the parametric inference process. We consider two distinct sample sizes, $m=\{200,500\}$, and set the maximum number of longitudinal responses to $d=10$. To replicate the unbalanced nature of the data, for each unit, we generate $n_{i}$ from a binomial distribution with $d$ trials and success probability $p=0.8$ (for $i=1,\dots,m$), thereby pruning the dataset. Given that the true parameters are known, we initially sample $\mathbf{U}_{i}$ (in a vectorized notation) from $n_{i}$-variate ($n_{i}\leq d$) $1$-factor and $2$-factor copulas (for $i=1,\dots,m$). Subsequently, we employ Probability Integral Transform (PIT) on $\mathbf{U}_{i}$ to generate per-unit responses $\mathbf{Y}_{i}$. Since the dimension for each response differs, we assume the same bivariate copulas for each factor with exchangeable parameters, resulting in a reduction in the number of estimable parameters from the model. We generate the continuous responses from the following model -
@@
where $\epsilon_{ij}(i.i.d)\sim N(0,1)$ (independent and identically distributed). Here we assign the regression coefficients as, $\beta_{0}=-0.5,\beta_{1}=-0.5,\beta_{2}=0.2,\beta_{3}=0.2$. Finally we generate the ordinal responses from the following model -